COVERING NUMBERS AND “LOW M^{∗}-ESTIMATE” FOR QUASI-CONVEX BODIES

A. E. LITVAK V. D. MILMAN A. PAJOR

Abstract. This article gives estimates on covering numbers and diameters of random proportional sections and projections of symmetric quasi-convex bodies in ℝ. These results were known for the convex case and played an essential role in development of the theory. Because duality relations can not be applied in the quasiconvex setting, new ingredients were introduced that give new understanding for the convex case as well.